OpenStudy (anonymous):

!!!! Need help at least starting it off!!! Pat and Jen volunteer at the library. Pat can shelve a full cart of books 14 minutes faster then Jen. If they work together, they can shelve a full cart in 24 min. How long does it take Jen to shelve a full cart by herself? ( helping me form a graph or someway to start will be extremely helpful.)

jimthompson5910 (jim_thompson5910):

Let P = time (in minutes) it takes for pat to do the job alone J = time (in minutes) it takes for jen to do the job alone we are told that " Pat can shelve a full cart of books 14 minutes faster then Jen", so that means Pat's time is 14 min less than Jen's time ie P = J - 14

jimthompson5910 (jim_thompson5910):

For Pat, it takes P minutes to do the full job That means he completes 1/P of the job in 1 minute You can think of it this way P minutes: 1 job 1 minute: 1/P job ... divided both parts of the ratio by P

jimthompson5910 (jim_thompson5910):

So Pat's rate is 1/P of a job per minute similarly, Jen's rate is 1/J of a job per minute ------------------------------------------------------- combine the two rates 1/P + 1/J J/(PJ) + P/(PJ) ( J + P )/(P*J) this combined rate is equal to 1/24. You can see this with this ratio 1 job: 24 minutes (for both working together) 1/24 job: 1 minute (divide both parts by 24) ------------------------------------------------------- So we have this equation ( J + P )/(P*J) = 1/24 cross multiplying gives you 24(J + P) = P*J now plug in P = J - 14 24(J + J - 14) = (J - 14)*J 24(J + J - 14) = J(J - 14) 24(2J - 14) = J(J - 14) 48J - 336 = J^2 - 14J I'll let you solve from here

OpenStudy (anonymous):

Alright I will start on it now, thank you for the start.